# Tree Ensembles for Contextual Bandits

### v1

@@ 4 Experiments
In this section, we evaluate TEUCB and TETS, and compare them against some of the best-known algorithms for solving the contextual bandit problem. For the implementations, we use the XGBoost library (Chen and Guestrin, 2016) to build gradient-boosted decision trees as our tree ensembles. We first evaluate the algorithms on a number of benchmark datasets. We then study them in the combinatorial contextual setting for solving the real-world problem of navigation over stochastic road networks in Luxembourg. The agent’s cumulative regret is calculated by comparing the path to an oracle agent who knows the expected travel duration for each edge given the current time of day. In practice, the oracle navigates the graph with zero regret. All the algorithms are evaluated w.r.t. cumulative regret averaged over ten random seeds.
@@ 4.1 Setup
Following the same method as Zhang et al. (2021), we use a simple procedure to transform a classification task into a contextual MAB problem. In each time step, the environment provides a context vector belonging to an unknown class. For the agent, each class corresponds to an arm, and the goal is to select the arm corresponding to the correct class. This results in a reward of 1, while an incorrect prediction results in a reward of 0. In practice, we note that our implementations of TEUCB and TETS seem to learn faster when their received rewards used for training are perturbed with a small amount of noise drawn from $\mathcal{N}(0,0.001)$. Though we see no apparent effect on the other models, we add the same sampled noise to the recorded rewards of every agent in the same time steps, for fairness.
@@
In the experiments in Section 4.2, we use the disjoint model for LinUCB, NeuralUCB, and NeuralTS. However, due to the nature of discrete splitting by decision trees, we can encode the context more effectively using a hybrid model for TEUCB and TETS. Therefore, we simply append the corresponding labels as a single character for each arm respectively at the beginning of the context: $\mathbf{x}_{k}=(k,\mathbf{x})$. Furthermore, we configure XGBoost to use partition-based splits (XGBoost Developers, 2022b) to ensure that TETS and TEUCB interpret the labels, and other categorical features in $\mathbf{x}$, as categorical rather than numerical. For LinUCB, NeuralUCB, and NeuralTS, categorical features are instead one-hot-encoded.
In these experiments, we set the time horizon to 10,000 time steps (except for the Mushroom dataset, where a horizon of 8,124 is sufficient to observe it entirely). For each of those time steps, the agents are presented with a feature vector $\mathbf{x}$ drawn randomly without replacement, which they encode for the different arms as described above. Subsequently, the agents predict the rewards for the individual arms according to the algorithms used. TEUCB (Algorithm 1) and TETS (Algorithm 2) are described in Section 3. The NeuralTS (Zhang et al., 2021), NeuralUCB (Zhou et al., 2020), and LinUCB (Li et al., 2010) algorithms are implemented according to their respective references.
@@ 4.2 Contextual Bandits
For this experiment, the data is collected from the UCI machine learning repository (Kelly et al., n.d.), and an overview of the four selected datasets is given in Table 1. It should be noted that Magic (Bock, 2007), which is short for Magic Gamma Telescope, contains only numerical features. The same is true for Shuttle (Statlog (n.d.), Shuttle), which is also known as Statlog. All features of Mushroom (Mushroom, 1987) are categorical, and Adult (Becker and Kohavi, 1996) has a balanced distribution between numerical and categorical features.
@@
Figure 1: Comparison of contextual MAB algorithms on UCI datasets.
@@ 4.2.1 Implementation
For all datasets in Table 1, the neural network agents use a network architecture of one hidden layer with 100 neurons. Regarding NeuralTS, NeuralUCB, and LinUCB, we search the same sets of hyperparameters as described by Zhang et al. (2021), who run these agents on the same datasets. One difference in our experiments is that we add a dropout probability of 0.2 when training the neural networks, which seems to have a positive impact on the performance. Furthermore, we let each of the networks train for 10 epochs.
For our TE bandits, we use XGBoost regressors with MSE loss and ensembles of 100 trees, as a way of matching the model complexity of the NeuralUCB and NeuralTS agents, and they are refitted to all previous observations in every time step. Our initial investigations have revealed that they are robust w.r.t. different choices of hyperparameters for XGBoost (XGBoost Developers, 2022c), and consequently, we settle on the default values. The only exceptions are the maximum tree depth, which we set to $10$, and the learning rate, which we set to $\eta=0.05$. For the additional parameters of TEUCB and TETS, we select $\nu=1$ and $T_{I}=10$.
The results are obtained using an NVIDIA A40 GPU for NeuralUCB and NeuralTS, and a desktop CPU for the other agents. To show the differences in computational performance between the agents, we report the runtime of the neural agents on the same CPU as well.
4.2.2 Experimental Results
The results of the experiments described above are illustrated in Fig. 1 and Table 2. We observe that both TEUCB and TETS consistently outperform all other models by a large margin and yield a significantly lower cumulative regret, where TEUCB slightly outperforms TETS. In addition to regret minimization, the CPU experiments indicate that TEUCB and TETS are significantly more efficient than their neural counterparts from a computational perspective. Notably, LinUCB is the most efficient method in terms of computations, however, it fails to adequately minimize the regret for our datasets.
@@ 4.2.2 Experimental Results
TETS
Mean 1548.9 1709.4 65.2 146.4
SD $\pm$ 40.2 $\pm$ 27.0 $\pm$ 13.1 $\pm$ 51.1
CPU 7 7 2 2
@@ 4.3 Combinatorial Contextual Bandits
In this section, we investigate the combinatorial contextual bandit methods on a real-world application, where we study two scenarios corresponding to performing the most efficient navigation over the real-world road network of Luxembourg. Similar navigation problems have recently been studied from a CMAB perspective, but often without contextual information (Åkerblom et al., 2023) or limited to neural bandit methods (Hoseini et al., 2022). CMAB methods are well-suited to the navigation problem, since the traversal time of each individual road segment can be highly stochastic and dependent on local factors (e.g., road works, traffic congestion, stop lights) about which knowledge may be gathered through sequential interactions with the environment.
Figure 2: Experimental results on real-world road network navigation in Luxembourg.
@@ 4.3.1 Implementation
We model the road network of Luxembourg via a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, with $|\mathcal{V}|=2,247$ vertices and $|\mathcal{E}|=5,651$ edges. The vertices in the graph represent intersections in the road network, and the edges represent individual road segments connecting the intersections. In this scenario, edges correspond to base arms, and paths (i.e., ordered sequences of edges) correspond to super arms. Each vertex has a coordinate consisting of longitude, latitude, and altitude values. For all edges $e\in\mathcal{E}$, the contextual vector $\mathbf{x}_{e}$ describes each road segment in the network. The agent is presented with a vector containing contextual data according to Table 3.
@@
$x$
Start position along x-axis.
$y$
Start position along y-axis.
$z$
Start position along z-axis.
$x^{\prime}$
End position along x-axis.
$y^{\prime}$
End position along y-axis
$z^{\prime}$
End position along z-axis
$\sqrt{(x-x^{\prime})^{2}}$
Euclidean distance along x-axis.
$\sqrt{(y-y^{\prime})^{2}}$
Euclidean distance along y-axis.
$\sqrt{(z-z^{\prime})^{2}}$
Euclidean distance along z-axis.
speed_limit
Maximum speed limit
stop
A boolean if edge includes a stop
time
The current time of day
Edge traversal times have been collected using the Luxembourg SUMO Traffic (LuST) simulation scenario (Codeca et al., 2015). The recorded edge traversal times are used to form kernel density estimators (KDE) (Weglarczyk, 2018) for each edge. If an edge does not contain any recorded traversals, the expected traversal time is set to the length of the edge divided by the vehicle speed. At each time step $t$, the time of day is randomly sampled, and used for updating the expected travel times of all edges and the corresponding KDEs. We generate edge-specific feedback by individually sampling travel times from the KDE of each edge on the chosen path.
As the graph contains more than 5,000 edges, the agents need to learn how the contextual features relate to the expected travel time. The road types in the graph are highways, arterial roads, and residential streets, with a total length of 955 km. In our experiments, we specifically study two problem instances (characterized by start and end nodes, as described in Fig. 2 and visualized in Appendix A). The agent predicts the expected travel time for each of the edges in the graph, and then solves the shortest path problem using Dijkstra’s algorithm (Dijkstra, 1959). The agents are evaluated on the sum of the expected travel times for all edges forming the traversed path compared to the expected travel time of the optimal path $\mathcal{S}^{*}$.
For this experiment, the neural agents utilize a neural network with two hidden layers where both hidden layers have 100 neurons, a dropout probability of 0.2, $\lambda=0.1$, $\nu=0.001$, and as in previous experiments, the network is trained over 10 epochs with the option of early stopping. For TETS and TEUCB, we utilize XGBoost regressors with default hyperparameters. The number of trees is $N=100$, the initial random selection is set to $T_{I}=10$, with the exploration factor $\nu=1$.
@@ 4.3.2 Experimental Results
Fig. 2 shows the results of TEUCB and TETS along with the baseline methods. We observe that TEUCB and TETS significantly outperform the other methods, both in how rapidly they learn and the steepness of the cumulative regret as $t$ increases. The neural methods both outperform LinUCB. In addition, our methods are much less computationally expensive than the neural methods. TETS, TEUCB and LinUCB were run on desktop CPUs while NeuralUCB and NeuralTS were run on NVIDIA V100 GPUs.
Both NeuralTS and NeuralUCB require large networks and substantial hyperparameter tuning to achieve deliberate exploration. In contrast, TETS and TEUCB achieve solid results as long as the exploration bonus is sufficiently small, such that edge weights have positive values. It is also noticeable that the neural agents are sensitive to the weight initialization of the networks, which leads to higher variance, see Fig. 2. The agents’ traversals are presented in Fig. 3 and Fig. 4, reported in Appendix A.
@@ 5 Conclusion
We developed a novel framework for contextual multi-armed bandits using tree ensembles. Within this framework, we incorporated the two commonly used methods: UCB and Thompson Sampling. Furthermore, we extended the framework to handle combinatorial contextual bandits, enabling more complex action selection at each time step. To demonstrate the effectiveness of the framework, we conducted extensive experiments on benchmark datasets, and utilized it for navigation over stochastic real-world road networks.

### v2

@@ 4 Experiments
In this section, we evaluate TEUCB and TETS, and compare them against several best-known algorithms for solving the contextual bandit problem. For the implementations, we use the XGBoost library (Chen and Guestrin, 2016) to build gradient-boosted decision trees as our tree ensembles, as well as the version of random forest found in the scikit-learn library (Pedregosa et al., 2011). We first evaluate the algorithms on benchmark datasets. We then study them in the combinatorial contextual setting for solving the real-world problem of navigation over stochastic road networks in Luxembourg. The agent’s cumulative regret is calculated by comparing the path to an oracle agent, who knows the expected travel duration for each edge given the current time of day. All the algorithms are evaluated w.r.t. cumulative regret averaged over ten random seeds and we report the average results.
@@ 4.1 Setup
We use the same procedure as Zhang et al. (2021) to transform a classification task into a contextual MAB problem. In each time step, the environment provides a context vector belonging to an unknown class. For the agent, each class corresponds to an arm, and the goal is to select the arm corresponding to the correct class. This results in a reward of 1, while an incorrect prediction results in a reward of 0.
@@
In the experiments in Section 4.2, we use the disjoint model for LinUCB, LinTS, NeuralUCB, and NeuralTS. However, due to the nature of discrete splitting by decision trees, we can encode the context more effectively using a hybrid model for tree-based methods. There, we simply append the corresponding labels as a single character for each arm respectively at the beginning of the context: $\mathbf{x}_{k}=(k,\mathbf{x})$, and mix numeric and categorical features in the context vectors. For LinUCB, LinTS, NeuralUCB, and NeuralTS, categorical features are one-hot-encoded.
In these experiments, we set the time horizon to 10,000 time steps (except for the Mushroom dataset, where a horizon of 8,124 is sufficient to observe it entirely). For each of those time steps, the agents are presented with a feature vector $\mathbf{x}$ drawn randomly without replacement, which they encode for the different arms as described above. Subsequently, the agents predict the rewards for the individual arms according to the algorithms used. TEUCB (Algorithm 1) and TETS (Algorithm 2) are described in Section 3. The TreeBootstrap (Elmachtoub et al., 2017), NeuralUCB (Zhou et al., 2020), NeuralTS (Zhang et al., 2021), LinUCB (Li et al., 2010) and LinTS (Agrawal and Goyal, 2013) algorithms are implemented according to their respective references.
@@ 4.2 Contextual Bandits
For this experiment, the data is collected from the UCI machine learning repository (Kelly et al., n.d.), where an overview of the used selected datasets is given in Table 1. It should be noted that Magic (Bock, 2007), which is short for Magic Gamma Telescope, contains only numerical features. The same is true for Shuttle (Statlog (n.d.), Shuttle), which is also known as Statlog. All features of Mushroom (Mushroom, 1987) are categorical, and Adult (Becker and Kohavi, 1996) has a balanced distribution between numerical and categorical features.
@@
Figure 1: Comparison of contextual MAB algorithms on UCI datasets. Figures 1(a), 1(b), 1(c), and 1(d) share the same color scheme for consistency, but the legend is only presented in 1(a) for improved visibility.
@@ 4.2.1 Implementation
For all datasets in Table 1, the neural network agents use a network architecture of one hidden layer with 100 neurons. Regarding NeuralUCB, NeuralTS, LinUCB, and LinTS, we search the same sets of hyper-parameters as described by Zhang et al. (2021), who run these agents on the same datasets, and select the parameters with the best performance. One difference in our experiments is that we add a dropout probability of 0.2 when training the neural networks, which seems to have a positive impact on the performance. Furthermore, we let each of the networks train for 10 epochs.
For all tree ensemble bandits, we use XGBoost and random forest regressors with MSE loss and ensembles of 100 trees.
Initial test runs revealed that the agents are robust w.r.t. different choices of hyper-parameters for XGBoost (XGBoost Developers, 2022b), and consequently, we settle on the default values. The only exception is the maximum tree depth, which we set to $10$. The same maximum tree depth is used for the implementation of random forests. For the exploration factor of TEUCB and TETS, we select $\nu=1$. As for the decision tree (DT) algorithm used for TreeBootstrap, we use the scikit-learn library to employ CART (Breiman et al., 2017), which is free of tunable parameters. This is similar to the implementation presented in the original work on TreeBootstrap (Elmachtoub et al., 2017).
Another point we noted during our initial test runs was that accurate predictions seem to be more important than less biased estimates obtained by splitting the previously observed samples into distinct data sets for ensemble fitting and value calculations respectively. Therefore, we use all observed context-reward pairs for both purposes in all experiments with TEUCB and TETS.
In order to avoid unnecessary computations and speed up the runs, TEUCB and TETS do not build new tree ensembles from scratch at each time step. Instead, they only consider which leaves the latest observation assigned in each tree and update the corresponding $o$, $s^{2}$, and $c$ parameters. Initially, when new observed data points may have a relatively large effect on the optimal tree ensemble, re-building happens more frequently. However, as the effect that the data points are expected to have on the tree architectures degrades over time, less re-building takes place. More precisely, re-building happens when the function $\lceil 8\ln(t)\rceil$ increases by one compared to the previous time step.
The experimental results are obtained using an NVIDIA A40 GPU for NeuralUCB and NeuralTS, and a desktop CPU for the other agents. To illustrate the differences in computational performance between the agents, we report the runtime of the neural agents on the same CPU as well.
@@
TETS-XGBoost
Mean 1548.0 1703.2 78.6 165.7
SD $\pm$ 30.9 $\pm$ 25.2 $\pm$ 11.8 $\pm$ 31.2
CPU 0.13 0.14 0.09 0.08
GPU - - - -
TEUCB-RF
Mean 1550.6 1678.5 58.2 190.8
SD $\pm$ 32.9 $\pm$ 31.3 $\pm$ 10.9 $\pm$ 68.9
CPU 0.12 0.13 0.10 0.12
GPU - - - -
TETS-RF
Mean 1566.3 1688.6 57.7 166.3
SD $\pm$ 34.2 $\pm$ 38.3 $\pm$ 8.1 $\pm$ 37.1
CPU 0.12 0.13 0.10 0.12
@@
LinTS
Mean 2253.2 2703.8 892.2 1263.0
SD $\pm$ 37.7 $\pm$ 30.4 $\pm$ 18.6 $\pm$ 22.1
CPU 0.02 0.02 0.04 0.11
GPU - - - -
TreeBootstrap-XGBoost
Mean 1693.2 1836.1 91.6 164.1
SD $\pm$ 18.8 $\pm$ 36.7 $\pm$ 5.6 $\pm$ 24.4
CPU 0.97 1.2 0.10 0.27
GPU - - - -
TreeBootstrap-RF
Mean 1584.3 1722.9 82.5 161.9
SD $\pm$ 32.5 $\pm$ 28.0 $\pm$ 3.3 $\pm$ 29.2
CPU 3.8 4.2 1.8 4.2
GPU - - - -
TreeBootstrap-DT
Mean 2187.2 2473.3 168.9 216.0
SD $\pm$ 57.1 $\pm$ 35.3 $\pm$ 11.0 $\pm$ 22.9
CPU 0.08 0.09 0.04 0.08
GPU - - - -
4.2.2 Experimental Results
The results of the experiments described above are presented in Fig. 1 and Table 2. We observe that the tree ensemble methods consistently outperform all other models by a large margin, both with XGBoost and random forest, and yield significantly lower cumulative regrets. Furthermore, TEUCB and TETS tend to perform better than TreeBootstrap on the adult, magic, and mushroom data sets, all of which contain only two arms.
On the shuttle data set, with seven different arms, the TreeBootstrap agents show comparable and even slightly better average performance in terms of regret minimization. It appears that TreeBootstrap’s assigning of a separate tree-based model for each arm is beneficial for this problem. However, this comes at the expense of having to fit multiple models in each time step, which is time-consuming and computationally demanding. Furthermore, its inability to generalize the reward predictions of distinct but related arms may in some cases be disadvantageous, especially for larger arm sets. Comparing TEUCB vs. TETS, and XGBoost vs. random forest, there is no clear winner as they all tend to perform comparable (and very effective) on different data sets. In addition to regret minimization, the CPU experiments indicate that TEUCB and TETS are significantly more efficient than their neural counterparts from a computational perspective. Notably, LinUCB, LinTS, and TreeBootstrap-DT are the most efficient methods in terms of computation. However, they fail to match the level of regret minimization as most other agents for our data sets.
@@ 4.3 Combinatorial Contextual Bandits
In this section, we investigate the combinatorial contextual bandit methods on a real-world application, where we study two scenarios corresponding to performing the most efficient navigation over the real-world road network of Luxembourg. This problem is crucial with the emergence of electric vehicles to mitigate the so-called range anxiety. Similar navigation problems have recently been studied from a CMAB perspective, but often without contextual information (Åkerblom et al., 2023) or limited to neural bandit methods (Hoseini et al., 2022). CMAB methods are well-suited to the navigation problem since the traversal time of each road segment can be highly stochastic and dependent on local factors (e.g., road works, traffic congestion, stop lights) about which knowledge may be gathered through sequential interactions with the environment.
Figure 2: Experimental results on real-world road network navigation in Luxembourg. Fig. 2(b) shares the legend provided in Fig. 2(a).
@@ 4.3.1 Implementation
We model the road network of Luxembourg via graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, with $|\mathcal{V}|=2,247$ vertices and $|\mathcal{E}|=5,651$ edges. The vertices represent intersections in the road network, and the edges represent individual road segments connecting the intersections. In this scenario, edges correspond to base arms, and paths (i.e., ordered sequences of edges) correspond to super arms. Each vertex has a coordinate consisting of longitude, latitude, and altitude values. For all edges $e\in\mathcal{E}$, the contextual vector $\mathbf{x}_{e}$ describes each road segment in the network. The agent is presented with a vector containing contextual data according to Table 3.
@@
$x$ Start position along x-axis.
$y$ Start position along y-axis.
$z$ Start position along z-axis.
$x^{\prime}$ End position along x-axis.
$y^{\prime}$ End position along y-axis
$z^{\prime}$ End position along z-axis
$\sqrt{(x-x^{\prime})^{2}}$ Euclidean distance along x-axis.
$\sqrt{(y-y^{\prime})^{2}}$ Euclidean distance along y-axis.
$\sqrt{(z-z^{\prime})^{2}}$ Euclidean distance along z-axis.
speed_limit Maximum speed limit
stop A boolean if edge includes a stop
time The current time of day
Edge traversal times have been collected using the Luxembourg SUMO Traffic (LuST) simulation scenario (Codeca et al., 2015). The recorded edge traversal times are used to form kernel density estimators (KDE) (Weglarczyk, 2018) for each edge. If an edge does not contain any recorded traversals, the expected traversal time is set to the length of the edge divided by the speed limit. At each time step $t$, the time of day is randomly sampled, and used for updating the expected travel times of all edges and the corresponding KDE’s. We generate edge-specific feedback by individually sampling travel times from the KDE of each edge on the chosen path.
As the graph contains more than 5,000 edges, the agents need to learn how the contextual features impact the expected travel time. The road types in the graph are highways, arterial roads, and residential streets, with a total length of 955 km. In our experiments, we specifically study two problem instances (characterized by different start and end nodes), referred to as problem instance 1 and problem instance 2. The paths selected by the different agents during a single run are visualized for problem instance 1 in Fig. 3.
An agent predicts the expected travel time for each of the edges in the graph, and then solves the shortest path problem using Dijkstra’s algorithm (Dijkstra, 1959). The agents are evaluated based on the sum of the expected travel times for all edges forming the traversed path compared to the expected travel time of the optimal path $\mathcal{S}^{*}$.
For this experiment, the neural agents utilize a neural network with two hidden layers, both containing 100 fully connected neurons. A dropout probability of 0.2 is used and after a parameter search over the same sets of values as in Section 4.2.1, we set $\lambda=0.1$ and $\nu=0.001$. Furthermore, the network is trained over 10 epochs with the option of early stopping if the MSE loss does not tend to keep improving. We implement the tree ensemble bandits utilizing XGBoost and random forest regressors with maximum tree depths of $10$; all other hyper-parameters being set to their default values. The number of trees is set to $N=100$ for all agents with tree ensembles, and the initial random selection is set to $T_{I}=10$. An exploration factor of $\nu=1$ is being used for TEUCB and TETS and the frequency of re-building their tree ensembles is the same as described in Section 4.2.1.
Figure 3: The road network of Luxembourg shows the trajectories of different agents for the experiment on problem instance 1, where the corresponding cumulative regret is presented in Fig. 2(a).
@@ 4.3.2 Experimental Results
Fig. 2 shows the results of the TEUCB and TETS methods along with the baselines. The road network and the agents’ traversals are presented in Fig. 3. The frequently an edge has been traversed is indicated by the red saturation, where more traversals correspond to a higher saturation. It is worth noting that LinUCB, LinTS, and the TreeBootstrap require excessive exploration as their models are edge-specific. On the contrary, TETS, TEUCB, NeuralTS, and NeuralUCB train one model with the possibility of generalizing the expected travel time predictions over different edges. We observe that using both XGBoost and random forest, TEUCB and TETS significantly outperform the other methods. In this setting, if an agent can generalize well what it has learned from one arm’s reward distribution to the other arms, it then can effectively avoid over-exploration, as indicated by both the regret plots in Fig. 2 and the agent’s specific path selections in Fig. 3. Comparing the two different tree ensemble methods, XGBoost seems to perform better on the navigation task compared to random forest in the TEUCB and TETS frameworks.
The neural methods tend to outperform LinUCB and LinTS. Compared to TreeBootstrap, however, the advantage is less clear, and seems to depend on the particular time horizon $T$. NeuralUCB and NeuralTS tend to learn faster, which is unsurprising since they can generalize reward predictions over different specific arms. After an initial period of heavy exploration, however, the TreeBootstrap agents appear to make good path selections more frequently, and yield cumulative reward curves that are more similar to those of TEUCB and TETS in slope.
All the agents were run on a single desktop CPU, apart from NeuralUCB and NeuralTS, which were run on an NVIDIA A40 GPU. NeuralTS and NeuralUCB tend to require large networks and substantial hyper-parameter tuning to achieve deliberate exploration, which can make them time-consuming while running and searching for good parameter values. In terms of run-time, the experiments took about an hour on the GPU for a neural agent. It is also noticeable that the neural agents appear to be sensitive to the weight initialization of the networks, which leads to higher variance (see Fig. 2). In contrast, TETS and TEUCB achieve solid results for a large range of parameter settings, yield smaller regret, and tend to run at comparable speed on a single CPU. The experiments with TEUCB and TETS took about 1.5 hours with XGBoost, and 5 hours using random forest as the tree ensemble method. The run-time of TreeBootstrap also tends to depend heavily on the particular tree model used. With a single decision tree per arm, the experiments took 0.5 hour, which is about the same as LinUCB and LinTS. Using tree ensembles, they took about 4 and 20 hours with XGBoost and random forests, respectively.
We note that TreeBootstrap with tree ensembles may have lower runtimes. Elmachtoub et al. (2017) propose a trick to speed up this method with a computational shortcut called TreeHeuristic. However, as acknowledged, its effectiveness is affected by some shortcomings including the inability to generalize over different but related arms through contextual information.
@@ 5 Conclusion
We developed a novel framework for contextual multi-armed bandits using tree ensembles. Within this framework, we incorporated the two commonly used methods for handling the exploration-exploitation dilemma: UCB and Thompson Sampling. Furthermore, we extended the framework to handle combinatorial contextual bandits, enabling more complex action selection at each time step.
To demonstrate the effectiveness of the framework, we conducted extensive experiments on benchmark datasets using the XGBoost and random forest algorithms. Additionally, we utilized it for navigation over stochastic real-world road networks.